If the product 3/2.4/3.5/4.6/5....a/b = 9. what is the sum of a and b?
Answer by wgomero(28)

Resolution:
3 4 5 6 18
- x - x - x - x...x - = 9
2 3 4 5 17
since this is a multiplication of fractions, you can cancel 3 on top and bottom, the same 4, 5 , 6, 7......17.
after canceling you the product end as follow:
18 = 9 where a = 18 and b = 2; the sum of a + b = 20.
-
2